# HQ-VAE: Hierarchical Discrete Representation Learning with Variational Bayes

### v1

@@ 4.3 Residual top-down layer
In this subsection, we set $R=1$ for demonstration purposes (the general case of $R$ is in Appendix C). This means the bottom-up and top-down paths are connected only at the top layer. We design a residual top-down layer for the approximated posterior, as in Figure 1. This layer is to better approximate the target feature with additional assignments of code vectors. By stacking this procedure $L$ times, the feature is approximated as
@@ 4.3.1 RSQ-VAE
where the numerator of the third term corresponds to the evaluation of the residuals $\bm{H}_{\bm{\phi}}(\bm{x})-\sum_{l^{\prime}=1}^{l}\bm{Z}_{l^{\prime}}$ for all $l\in[L]$ with the dequantization process. However, we empirically found that training the model with this ELBO objective above was often unstable. We suspect this is because the objective regularizes $\bm{Z}_{1:L}$ to make $\sum_{l^{\prime}=1}^{l}\bm{Z}_{l^{\prime}}$ close to the feature for all $l\in[L]$. We hypothesize that the regularization is too strong to regularize the latent representation. To address the issue, we will impose conditional distributions not on $(\bm{Z}_{1:L},\tilde{\bm{Z}}_{1:L})$ but on $(\bm{Z}_{1:L},\tilde{\bm{Z}})$, where $\tilde{\bm{Z}}=\sum_{l=1}^{L}\tilde{\bm{Z}}_{l}$. From the reproductive property of the Gaussian distribution, a continuous latent variable converted from $\bm{Z}$ via the stochastic dequantization processes, $\tilde{\bm{Z}}=\sum_{l=1}^{L}\tilde{\bm{Z}}_{l}$, follows a Gaussian distribution:
@@
$\mathcal{J}_{\text{RSQ-VAE}}=\frac{D}{2}\log\sigma^{2}+\mathbb{E}_{\mathcal{Q}(\bm{Z}_{1:L},\tilde{\bm{Z}}_{1:L}|\bm{x})}\left[\frac{\|\bm{x}-\bm{f}_{\bm{\theta}}(\bm{Z}_{1:L})\|_{2}^{2}}{2\sigma^{2}}+\frac{\|\tilde{\bm{Z}}-\bm{Z}\|_{F}^{2}}{2\sum_{l=1}^{L}s_{l}^{2}}-\sum_{l=1}^{L}H(\hat{P}_{s_{l}^{2}}(\bm{Z}_{l}|\tilde{\bm{Z}}_{l}))\right]$ (14)
@@
Remark. HQ-VAE has favorable properties similar to those of SQ-VAE. The training scheme does not require hyperparameters except for the temperature parameter of the Gumbel-softmax approximation (see Equations (6) and (14)). Furthermore, the derived models can benefit from the self-annealing effect, as in SQ-VAE; this is empirically shown in Section 5.3.
(a) CIFAR10
@@ 5 Experiments
We comprehensively examine SQ-VAE-2 and RSQ-VAE and their applicability to generative modeling. In particular, we compare SQ-VAE-2 with VQ-VAE-2 and RSQ-VAE with RQ-VAE to see if our framework improves reconstruction performance relative to the baselines. We basically compare various latent capacities in order to evaluate our methods in a rate–distortion (RD) sense (Alemi et al., 2017; Williams et al., 2020). The error plots the RD curves in the figures below indicate standard deviations based on four runs with different training seeds. In addition, we test HQ-VAE on an audio dataset to see if it is applicable to a different modality. Moreover, we investigate the characteristics of the injected top-down and residual top-down layers through visualizations. Lastly, we apply HQ-VAEs to generative tasks to demonstrate their feasibility as feature extractors. Unless otherwise noted in what follows, we use the same network architecture in all models and set the codebook dimension to $d_{b}=64$. The details of these experiments are in Appendix D.
@@
(a) Reconstructed images and magnified differences of SQ-VAE-2
(b) $H(\hat{P}_{s_{l}^{2}}(\bm{z}_{l,i}|\tilde{\bm{Z}}_{l}))$ in SQ-VAE-2
(c) Reconstructed images and magnified differences of RSQ-VAE
@@
Figure 4: Reconstructed samples with partial layers in (a) SQ-VAE-2 and (c) RSQ-VAE. The top row shows the reconstructed images, while the bottom row shows the components added at each layer. $(d_{l},K_{l})=(16^{2},256),(32^{2},16),(64^{2},4)$ and $(d_{l},K_{l})=(32^{2},4),(32^{2},16),(32^{2},256)$ for latent capacities, $l$ of 1, 2, and 3, respectively. Notice that the numbers of bits in these models are equal at each layer. For a reasonable visualization, we apply progressive coding, which induces progressive compression, to SQ-VAE-2 (see Appendix D.6). (b) and (d) Variance parameter $s_{l}^{2}$ normalized by the initial value $s_{l,0}^{2}$ and average entropy of the quantization process ($H(\hat{P}_{s_{l}^{2}}(\bm{z}_{l,i}|\tilde{\bm{Z}}_{l}))$) at each layer.
Validation on an audio dataset. We validate RSQ-VAE in the audio domain by comparing it with RQ-VAE in an experiment on reconstructing the normalized log-Mel spectrogram in an environmental sound dataset, UrbanSound8K (Salamon et al., 2014). We use the same network architecture as in an audio generation paper (Liu et al., 2021), in which multi-scale convolutional layers of varying kernel sizes are deployed to capture the local and global features of audio signals in the time-frequency domain (Xian et al., 2021). The codebook size is set to $K_{l}=8$. The number of layers is set to 4 and 8, and all the layers share the same codebook. We run each trial with five different random seeds and obtain the average and standard deviation of the RMSEs. As shown in Table 2, RSQ-VAE has better average RMSEs than RQ-VAE has across different numbers of layers on the audio dataset. We also evaluate the results in terms of perceptual quality by performing subjective listening tests, which are described in Appendix D.2.3.
@@ 5.3 Empirical study of top-down layers
In Appendix D.3, we explore the idea of combining the two layers to form a hybrid model. There it is shown that the individual layers in a hybrid model produce effects similar to using them alone. That is, the outputs from the injected top-down layers have better resolution and residual top-down layers make refinements upon certain decompositions. Since these two layers enjoy distinct refining mechanisms, a hybrid model may bring a more flexible approximation to the posterior distribution.
@@
${}^{\dagger}$ (ours) 9.74
${}^{\ddagger}$ (ours) 8.46
@@ 5.4 Applications of HQ-VAEs to generative tasks
First, we train RSQ-VAE on FFHQ (256$\times$256) by using the same encoder–decoder architecture as in Lee et al. (2022a). We set the latent capacities to $L=4$, $(d_{l},K_{l})=(8^{2},256)$ for $l\in[4]$ by following their RQ-VAE. We train prior models, an RQ-Transformer (Lee et al., 2022a) and a contextual RQ-Transformer (Lee et al., 2022b), on the latent feature extracted by RSQ-VAE. Table 4 lists the Fréchet inception distance (FID) scores for the generated images (refer to Table 6 for more detailed comparisons). According to the table, our RSQ-VAE leads to comparable generation performance to RQ-VAE, even without adversarial training.
Next, we train SQ-VAE-2 on ImageNet (256$\times$256) by using a simple resblock-based encoder–decoder architecture. We define two levels of discrete latent spaces whose capacities are $(d_{1},K_{1})=(16^{2},2048)$ and $(d_{2},K_{2})=(32^{2},1024)$. For training SQ-VAE-2 training, we use the adversarial training framework to achieve a perceptual reconstruction with a feasible compression ratio. Subsequently, we use Muse (Chang et al., 2023) to train the prior model on the hierarchical discrete latent features. As reported in Table 4, our SQ-VAE-2-based models achieve FID score that is competitive with current state-of-the-art models (refer to Table 7 for more detailed comparisons).
@@ 6.2 Concluding remarks
Appendix A From SQ-VAE to HQ-VAE
@@ B.2 RSQ-VAE
By substituting Equations (17) and (20) into (19), we find that Equation (14), where we use $\tilde{\bm{Z}}=\bm{H}_{\bm{\phi}}(\bm{x})$ instead of sampling it in a practical implementation. The above is the derivation of the ELBO objective in the case of $R=1$. Appendix C extends this model to the general case with a variable $R$, which is equivalent to the hybrid model.
Appendix C Hybrid model
Figure 6: Top-down layers corresponding to the $r$th resolution of the hybrid model in Appendix C.
@@ Appendix C Hybrid model
Appendix D Experimental details
This appendix describes the details of the experiments [^2] The source code is attached in the supplementary material. discussed in Section 5. For all the experiments except for the ones of RSQ-VAE and RQ-VAE on FFHQ and UrbanSound8K in Section 5.2, we construct architectures for the bottom-up and top-down paths as is described in Figures 1 and 7. To build these paths, we use two common blocks, the Resblock and Convblock by following Child (2021) in Figure 6(a); these blocks are shown in Figures 6(b) and 6(c). Here, we denote the width and height of $\bm{H}_{\bm{\phi}}^{r}(\bm{x})$ as $w_{r}$ and $h_{r}$, respectively, i.e., $\bm{H}_{\bm{\phi}}^{r}(\bm{x})\in\mathbb{R}^{d_{b}\times w_{r}\times h_{r}}$. We set $c_{\mathrm{mid}}=0.5$ in Figure 7. For all the experiments, we use the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.9$. Unless otherwise noted, we reduce the learning rate in half if the validation loss does not improve in the last three epochs.
In HQ-VAE, we deal with the decoder variance $\sigma^{2}$ by using the update scheme with the maximum likelihood estimation (Takida et al., 2022a). We gradually reduce the temperature parameter of the Gumbel–softmax trick with a standard schedule $\tau=\exp(10^{-5}\cdot t)$ (Jang et al., 2017), where $t$ is the iteration step.

### v2

@@ 4.3 Residual top-down layer
In this subsection, we set $R=1$ for demonstration purposes (the general case of $R$ is in Appendix D). This means the bottom-up and top-down paths are connected only at the top layer. In this setup, the bottom-up path generates only a single-resolution feature; thus, we denote $\bm{H}_{\bm{\phi}}^{1}(\bm{x})$ as $\bm{H}_{\bm{\phi}}(\bm{x})$ for notational simplicity in this subsection. We design a residual top-down layer for the approximated posterior, as in Figure 1. This layer is to better approximate the target feature with additional assignments of code vectors per encoded vector. By stacking this procedure $L$ times, the feature is approximated as
@@ 4.3.1 RSQ-VAE
where the numerator of the third term corresponds to the evaluation of the residuals $\bm{H}_{\bm{\phi}}(\bm{x})-\sum_{l^{\prime}=1}^{l}\bm{Z}_{l^{\prime}}$ for all $l\in[L]$ with the dequantization process. However, we empirically found that training the model with this ELBO objective above was often unstable. We suspect this is because the objective regularizes $\bm{Z}_{1:L}$ to make $\sum_{l^{\prime}=1}^{l}\bm{Z}_{l^{\prime}}$ close to the feature for all $l\in[L]$. We hypothesize that the regularization excessively pushes information to the top layers, which may result in layer collapse in the bottom layers. To address the issue, we modify the prior distribution $\mathcal{P}$ to result in weaker latent regularizations. Specifically, we define the prior distribution as a joint distribution of $(\bm{Z}_{1:L},\tilde{\bm{Z}})$ instead of that of $(\bm{Z}_{1:L},\tilde{\bm{Z}}_{1:L})$, where $\tilde{\bm{Z}}=\sum_{l=1}^{L}\tilde{\bm{Z}}_{l}$. From the reproductive property of the Gaussian distribution, a continuous latent variable converted from $\bm{Z}$ via the stochastic dequantization processes, $\tilde{\bm{Z}}=\sum_{l=1}^{L}\tilde{\bm{Z}}_{l}$, follows a Gaussian distribution:
@@
$\mathcal{J}_{\text{RSQ-VAE}}=\frac{D}{2}\log\sigma^{2}+\mathbb{E}_{\mathcal{Q}(\bm{Z}_{1:L},\tilde{\bm{Z}}_{1:L}|\bm{x})}\left[\frac{\|\bm{x}-\bm{f}_{\bm{\theta}}(\bm{Z}_{1:L})\|_{2}^{2}}{2\sigma^{2}}+\frac{\|\tilde{\bm{Z}}-\bm{Z}\|_{F}^{2}}{2\sum_{l=1}^{L}s_{l}^{2}}-\sum_{l=1}^{L}H(\hat{P}_{s_{l}^{2}}(\bm{Z}_{l}|\tilde{\bm{Z}}_{l}))\right]{\color[rgb]{0,0,0}{{,}}}$ (14)
@@
Remark. HQ-VAE has favorable properties similar to those of SQ-VAE. The training scheme significantly reduces the number of hyperparameters required for training VQ-VAE variants. First, the objective functions (6) and (14) do not include any hyperparameters such as $\beta$ in Equations (7) and (15), except for the introduction of a temperature parameter for the Gumbel-softmax to enable the reparameterization of the discrete latent variables (Jang et al., 2017; Maddison et al., 2017). Please refer to Appendix E for the approximation of the categorical distributions $\mathcal{Q}(\bm{Z}_{1:L},\tilde{\bm{Z}}_{1:L}|\bm{x})$. Moreover, our formulations do not require heuristic techniques such as stop-gradient, codebook reset, or EMA update. Additionally, HQ-VAE has several other empirical benefits over the VQ-VAE variants, such as the effect of self-annealing for better codebook utilization and a better rate-distortion (RD) trade-off, which will be verified in Section 5.
@@ 5 Experiments
We comprehensively examine SQ-VAE-2 and RSQ-VAE and their applicability to generative modeling. In particular, we compare SQ-VAE-2 with VQ-VAE-2 and RSQ-VAE with RQ-VAE to see if our framework improves reconstruction performance relative to the baselines. We basically compare various latent capacities in order to evaluate our methods in a RD sense (Alemi et al., 2017; Williams et al., 2020). The error plots the RD curves in the figures below indicate standard deviations based on four runs with different training seeds. In addition, we test HQ-VAE on an audio dataset to see if it is applicable to a different modality. Moreover, we investigate the characteristics of the injected top-down and residual top-down layers through visualizations. Lastly, we apply HQ-VAEs to generative tasks to demonstrate their feasibility as feature extractors. Unless otherwise noted in what follows, we use the same network architecture in all models and set the codebook dimension to $d_{b}=64$. Following the common dataset splits, we utilize the validation sets to adjust the learning rate during training and compute all the numerical metrics on the test sets. The details of these experiments are in Appendix F.
@@
Figure 4: Reconstructed samples with partial layers in (a) SQ-VAE-2 and (c) RSQ-VAE. The top row shows the reconstructed images, while the bottom row shows the components added at each layer. $(d_{l},K_{l})=(16^{2},256),(32^{2},16),(64^{2},4)$ and $(d_{l},K_{l})=(32^{2},4),(32^{2},16),(32^{2},256)$ for latent capacities, $l$ of 1, 2, and 3, respectively. Notice that the numbers of bits in these models are equal at each layer. For a reasonable visualization, we apply progressive coding, which induces progressive compression, to SQ-VAE-2 (see Appendix F.6). (b) and (d) Variance parameter $s_{l}^{2}$ normalized by the initial value $s_{l,0}^{2}$ and average entropy of the quantization process ($H(\hat{P}_{s_{l}^{2}}(\bm{z}_{l,i}|\tilde{\bm{Z}}_{l}))$) at each layer.
Validation on an audio dataset. We validate RSQ-VAE in the audio domain by comparing it with RQ-VAE in an experiment on reconstructing the normalized log-Mel spectrogram in an environmental sound dataset, UrbanSound8K (Salamon et al., 2014). We use the same network architecture as in an audio generation paper (Liu et al., 2021), in which multi-scale convolutional layers of varying kernel sizes are deployed to capture the local and global features of audio signals in the time-frequency domain (Xian et al., 2021). The codebook size is set to $K_{l}=8$. The number of layers is set to 4 and 8, and all the layers share the same codebook. We run each trial with five different random seeds and obtain the average and standard deviation of the RMSEs. As shown in Table 2, RSQ-VAE has better average RMSEs than RQ-VAE has across different numbers of layers on the audio dataset. We also evaluate the results in terms of perceptual quality by performing subjective listening tests, which are described in Appendix F.2.3.
@@ 5.3 Empirical study of top-down layers
In Appendix F.3, we explore the idea of combining the two layers to form a hybrid model. There it is shown that the individual layers in a hybrid model produce effects similar to using them alone. That is, the outputs from the injected top-down layers have better resolution and residual top-down layers make refinements upon certain decompositions. Since these two layers enjoy distinct refining mechanisms, a hybrid model may bring a more flexible approximation to the posterior distribution.
@@
RSQ-VAE (ours) 9.74
RSQ-VAE (ours) 8.46
Table 4: Image generation on ImageNet.
@@ 5.4 Applications of HQ-VAEs to generative tasks
First, we train RSQ-VAE on FFHQ (256$\times$256) by using the same encoder–decoder architecture as in Lee et al. (2022a). We set the latent capacities to $L=4$, $(d_{l},K_{l})=(8^{2},256)$ for $l\in[4]$ by following their RQ-VAE. We train prior models, an RQ-Transformer (Lee et al., 2022a) and a contextual RQ-Transformer (Lee et al., 2022b), on the latent feature extracted by RSQ-VAE. Table 4 lists the Fréchet inception distance (FID) scores for the generated images (refer to Table 7 for more detailed comparisons). According to the table, our RSQ-VAE leads to comparable generation performance to RQ-VAE, even without adversarial training. Figure 16 shows the generated samples from our model on FFHQ.
Next, we train SQ-VAE-2 on ImageNet (256$\times$256) by using a simple resblock-based encoder–decoder architecture. We define two levels of discrete latent spaces whose capacities are $(d_{1},K_{1})=(16^{2},2048)$ and $(d_{2},K_{2})=(32^{2},1024)$. For training SQ-VAE-2 training, we use the adversarial training framework to achieve a perceptual reconstruction with a feasible compression ratio. Subsequently, we use Muse (Chang et al., 2023) to train the prior model on the hierarchical discrete latent features. As reported in Table 4, our SQ-VAE-2-based models achieve FID score that is competitive with current state-of-the-art models (refer to Table 8 for more detailed comparisons). Figure 17 shows the generated images from our model on ImageNet.
@@ 6.2 Concluding remarks
Acknowledgement
We would like to thank Marc Ferras for many helpful comments during the discussion phase. Besides, we thank anonymous reviewers for their valuable suggestions and comments.
Appendix A Literature of stochastic posterior modeling for discrete latent
We briefly compare SQ-VAE, the foundation of HQ-VAE, with two other stochastic discrete models.
Sønderby et al. (2017) proposed stochastic posterior modeling in the discrete latent setting as $Q(\bm{z}_{i}=\bm{b}_{k}|\bm{x})\propto-\|\tilde{\bm{z}}_{i}-\bm{b}_{k}\|_{2}^{2}$, followed by succeeding works (Roy et al., 2018; Williams et al., 2020). Although the approach resembles SQ-VAE in terms of stochastic quantization, they have non-trivial differences as follows. In SQ-VAE, Takida et al. (2022b) introduced a trainable parameter, denoted as $\bm{\Sigma}_{\varphi}$, by considering a dequantization process as the inverse operation of quantization. This leads to $Q(\bm{z}_{i}=\bm{b}_{k}|\bm{x})\propto-\frac{1}{2}(\tilde{\bm{z}}_{i}-\bm{b}_{k})^{\top}\bm{\Sigma}_{\varphi}^{-1}(\tilde{\bm{z}}_{i}-\bm{b}_{k})$, which imposes the self-annealing effect as well as allows the posterior modeling with more general distances. Please refer to Takida et al. (2022b) for more details about the difference and some numerical examples.
The logits modeling (LM) in dVAE (Ramesh et al., 2021) is another approach to model the discrete posterior distribution ($Q(\mathbf{Z}|\mathbf{x})$) instead of vector quantization. LM assumes an encoder that directly maps from $\mathbf{x}$ to the logits of $Q(\mathbf{z}_{i}|\mathbf{x})$. Therefore, it does not have raw explicit features $\tilde{\mathbf{Z}}$. On the other hand, in SQ-VAE, the encoded features $\tilde{\mathbf{Z}}$ are directly quantized. When the learnable variance parameter $s^{2}$ is set (fixed) close to 00, conventional VQ is obtained. We used SQ instead of LM for the posterior modeling in HQ-VAE because SQ can serve as a drop-in replacement of VQ and can be extended to hierarchical and residual quantizations similar to VQ.
Appendix B From SQ-VAE to HQ-VAE
@@ C.2 RSQ-VAE
By substituting Equations (17) and (20) into (19), we find that Equation (14), where we use $\tilde{\bm{Z}}=\bm{H}_{\bm{\phi}}(\bm{x})$ instead of sampling it in a practical implementation. The above is the derivation of the ELBO objective in the case of $R=1$. Appendix D extends this model to the general case with a variable $R$, which is equivalent to the hybrid model.
Appendix D Hybrid model
Figure 6: Top-down layers corresponding to the $r$th resolution of the hybrid model in Appendix D.
@@ Appendix D Hybrid model
Appendix E Gumbel-softmax approximation
Since HQ-VAE has discrete latent space, our objective functions (6), (14), and (26) involve expectations with respective to the categorical variables $\bm{Z}_{1:L}$ (in $\mathcal{Q}(\bm{Z}_{1:L},\tilde{\bm{Z}}_{1:L}|\bm{x})$). As a common method to make the sampling process of the categorical variables reparameterizable, we replace the categorical distribution (2) with Gumbel-softmax alternatives. In a Gumbel-softmax distribution, a temperature parameter $\tau\in(0,\infty)$ is introduced. The sampling process from a Gumbel-softmax distribution is equivalent to that from a categorical distribution as $\tau\to 0$.
There is a trade-off with the temperature $\tau$. Lower temperatures lead to sampling that closely resembles the original categorical distribution but with larger gradient variances. Conversely, higher temperatures result in greater discrepancies between the categorical and Gumbel-softmax distributions but with smaller gradient variances. It has been shown that annealing the temperature from high to low values during model training performs well in practice (Jang et al., 2017; Sønderby et al., 2017; Williams et al., 2020). Specifically, Jang et al. (2017) proposed an exponential scheduling, $\tau=\max\{c_{\tau},\exp(-rt)\}$, where $t$ denotes the global training step. Following the approach of Takida et al. (2022b), we adopt the same annealing schedule with $c_{\tau}=0$ and $r=10^{-5}$ for all the experiments (except for Table 5).
Here, we investigate the sensitivity of the choice of the decay factor $r$. We train two-layer SQ-VAE-2 and RSQ-VAE with codebook capacities of $(d_{l},K_{l})=(4^{2},512),(8^{2},512)$ and $(d_{l},K_{l})=(8^{2},32),(8^{2},32)$, respectively. We sweep $r$ while keeping the other settings in Table 5. According to the table, a decay factor that is too large, leading to fast convergence to the categorical distribution, deteriorates the reconstruction performance.
Table 5: Impact of decay factor $r$ in the exponential schedule defined in Appendix E on reconstruction of images in CIFAR10. RMSE (×$\times 10^{2}$) is evaluated using the test set. Too large decay factor deteriorates the reconstruction performance.
Model $r=1\times 10^{-6}$ $r=3\times 10^{-6}$ $r=1\times 10^{-5}$ $r=3\times 10^{-5}$ $r=1\times 10^{-4}$ $r=3\times 10^{-4}$
SQ-VAE-2 5.223 $\pm$ 0.016 5.219 $\pm$ 0.017 5.264 $\pm$ 0.016 5.356 $\pm$ 0.016 7.234 $\pm$ 0.021 9.083 $\pm$ 0.019
RSQ-VAE 6.544 $\pm$ 0.018 6.149 $\pm$ 0.023 5.697 $\pm$ 0.022 5.579 $\pm$ 0.015 5.802 $\pm$ 0.020 6.439 $\pm$ 0.019
Appendix F Experimental details
This appendix describes the details of the experiments [^2] The source code is attached in the supplementary material. discussed in Section 5. For all the experiments except for the ones of RSQ-VAE and RQ-VAE on FFHQ and UrbanSound8K in Section 5.2, we construct architectures for the bottom-up and top-down paths as is described in Figures 1 and 7. To build these paths, we use two common blocks, the Resblock and Convblock by following Child (2021) in Figure 7(a); these blocks are shown in Figures 7(b) and 7(c). Here, we denote the width and height of $\bm{H}_{\bm{\phi}}^{r}(\bm{x})$ as $w_{r}$ and $h_{r}$, respectively, i.e., $\bm{H}_{\bm{\phi}}^{r}(\bm{x})\in\mathbb{R}^{d_{b}\times w_{r}\times h_{r}}$. We set $c_{\mathrm{mid}}=0.5$ in Figure 7. For all the experiments, we use the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.9$. Unless otherwise noted, we reduce the learning rate in half if the validation loss does not improve in the last three epochs.
In HQ-VAE, we deal with the decoder variance $\sigma^{2}$ by using the update scheme with the maximum likelihood estimation (Takida et al., 2022a). We gradually reduce the temperature parameter of the Gumbel-softmax trick with a standard schedule $\tau=\exp(10^{-5}\cdot t)$ (Jang et al., 2017), where $t$ is the iteration step.